# Deflection angle in the strong deflection limit: a perspective from local geometrical invariants and matter distributions

### v3

@@ I Introduction
Building on these insights, numerous studies have focused on developing analytical formulations for the deflection angle in the strong deflection limit and exploring their applications. An explicit expression for the deflection angle in this context was first derived in Ref. [5], establishing the foundation for subsequent research. Later, this approach was refined in Refs. [6, 7], thus extending its applicability to various spacetimes. Due to significant astrophysical interest in this regime and the methodological simplicity of these techniques, they have been applied to a broad range of compact astrophysical objects. Notable examples include the Schwarzschild black hole [8], the Reissner–Nordström black hole [9, 10], regular black holes [11], naked singularities [12, 13], wormholes [14, 15, 16, 17, 18, 19, 20], ultracompact objects [7], gravastars [21], and other black holes [22, 23]. Additionally, formulations in the strong deflection limit have been extended to incorporate finite-distance corrections [24, 25] and to describe the deflection of massive particles [26]. Consequently, the logarithmic divergence of the deflection angle in this regime has been consistently verified across a wide range of spacetimes.
@@
Here, $b$ denotes the impact parameter, and $b_{\mathrm{c}}$ is its critical value. This formula has been successfully applied to a wide variety of astrophysical scenarios. However, two fundamental issues remain unresolved in the current formulation. First, the physical interpretation of $\bar{a}$ and $\bar{b}$ is not yet fully understood. Although these coefficients are expressed in terms of the local values of metric functions and their derivatives (and are therefore practically useful), their precise relationship to physical quantities—such as the local spacetime geometry or the influence of matter fields—remains unclear. Second, these coefficients, in their current form, are coordinate-dependent. Since they are related to observable quantities, it is essential to reformulate them in a coordinate-independent manner in accordance with the principles of relativity, thus enabling a more universal physical interpretation.
@@ II Photon dynamics and circular orbits in static, spherically symmetric spacetimes
where $t$, $r$, and $(\theta,\varphi)$ denote the time, radial, and angular coordinates, respectively. The functions $\nu(r)$, $\lambda(r)$, and $R(r)$ are arbitrary; however, one of these can be fixed by an appropriate gauge choice. Here, $R(r)$ represents the areal radius (i.e., $R(r)>0$). We assume that the spacetime is asymptotically flat, meaning that as $R\to\infty$ (corresponding to $r\to r_{\infty}$, with $r_{\infty}$ denoting the coordinate value at spatial infinity), the metric functions approach $e^{\nu}\to 1$ and $e^{\lambda}\to(R^{\prime})^{2}$, ensuring convergence to the Minkowski metric, [^1] In our formulation, we do not impose the coordinate condition that the radial coordinate asymptotically approaches the areal radius at infinity, as is commonly done in the literature (e.g., Refs. [5, 6]). where, and hereafter, the prime denotes differentiation with respect to $r$.
@@
provided that $R>2m$, which ensures that the metric is well-defined.
@@
Due to the spherical symmetry of the metric, their orbits can be confined to the equatorial plane (i.e., $\theta=\pi/2$) without loss of generality. Under these assumptions, the Lagrangian for a photon is given by
@@ III Derivation of the deflection angle in the strong deflection limit
Following Ref. [5], we evaluate the deflection angle in the strong deflection limit. We introduce the new variable
@@
which differs from the variable used in the literature. This definition ensures that the final expressions are formulated primarily in terms of the areal radius. As a result, the formulation is both coordinate-invariant and geometrically meaningful. With this substitution, the integral $I(r_{0})$ becomes
@@ III.1 $R^{\prime}_{\mathrm{m}}\neq 0$
$c_{2}$ $=-\frac{V^{\prime\prime}_{\mathrm{m}}}{2}+O\left(\frac{r_{0}}{r_{\mathrm{m}}}-1\right),$ (30)
@@
While previous works in the literature have expressed these coefficients arising from $I_{\mathrm{D}}$ in terms of metric functions, our formulation instead recasts them in terms of local quantities at the photon sphere (i.e., $V^{\prime\prime}_{\mathrm{m}}$ and $R_{\mathrm{m}}$) together with a well-defined quasi-local quantity (i.e., $m(r_{\mathrm{m}})$). This approach not only clarifies the physical origin of the logarithmic divergence in the deflection angle but also identifies the observable quantities governing the phenomenon in the strong deflection limit.
@@ III.2 $R^{\prime}_{\mathrm{m}}=0$
Next, we assume that $R^{\prime}_{\mathrm{m}}=0$, which implies that the spacetime exhibits a wormhole geometry with the photon sphere located at the throat. In this case, the expansion of $c_{1}$ and $c_{2}$ around $r_{0}=r_{\mathrm{m}}$ reduces to [^3] Although the first-order term in $(r_{0}/r_{\mathrm{m}}-1)$ in the expansion of $c_{2}$ does not affect the strong field limit coefficients, we retain it to capture the unique behavior of the higher-order correction terms for the Ellis wormhole.
@@ IV Formula for the deflection angle in the strong deflection limit
$\nu^{\prime}$ $=\frac{Re^{\lambda}}{R^{\prime}}G_{(1)(1)}-\frac{(R^{\prime})^{2}-e^{\lambda}}{RR^{\prime}},$ (53)
@@
Consequently, these results imply that information about the curvature near the photon sphere can be extracted from measurements of the deflection angle in the strong deflection regime. Specifically, if $\bar{a}$ and $R_{\mathrm{m}}$ are determined from observations, the curvature combination $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$ can be directly derived. Moreover, if $\bar{b}$ is also measured and a specific spacetime model is assumed—so that all the relevant quantities can be computed for direct comparison with observations—then we can deduce $G_{(1)(1)}^{\mathrm{m}}$ (with $m(r_{\mathrm{m}})$) determined via Eq. (55) in the case $R^{\prime}_{\mathrm{m}}\neq 0$, or $G_{(0)(0)}^{\mathrm{m}}$ (with $\mathrm{d}m/\mathrm{d}R|_{r=r_{\mathrm{m}}}$) determined via Eq. (56) in the case $R^{\prime}_{\mathrm{m}}=0$. In other words, if the theoretical predictions for $\bar{a}$ and $\bar{b}$ agree reasonably well with the observed values, this approach enables us to infer the local curvature and the gravitational mass, thereby providing a robust consistency check of the underlying gravitational model.
@@ V Interpreting strong deflection in terms of local matter fields in general relativity
Finally, we express the strong field limit coefficients $\bar{a}$ and $\bar{b}$ in terms of these matter field quantities. By substituting Eqs. (62) into Eqs. (60) and (69), we obtain
@@
As the simplest case, we consider a vacuum spacetime in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Schwarzschild line element
@@
Next, we consider an electrovacuum spacetime (i.e., a vacuum spacetime containing only electromagnetic fields) in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Reissner–Nordström line element
@@
Next, consider a special case in which the spacetime is nonvacuum, but matter fields satisfy $\rho(r_{\mathrm{m}})+\Pi(r_{\mathrm{m}})=0$. In this situation, the strong field limit coefficient $\bar{a}$ remains equal to $1$, thereby mirroring the Schwarzschild result despite the presence of nonvanishing matter fields. This demonstrates that when $\rho(r_{\mathrm{m}})+\Pi(r_{\mathrm{m}})=0$, the strong deflection limit exhibits universal behavior with $\bar{a}=1$, irrespective of the specific details of the matter content.
@@
which agree with those in Ref. [14], employing the same expansion parameter used here. Note that for the Ellis wormhole under consideration, the conditions $R^{\prime\prime\prime}_{\mathrm{m}}=0$ and $V^{\prime\prime\prime}_{\mathrm{m}}=0$ are satisfied, which causes the coefficient of $(b/b_{\mathrm{c}}-1)^{1/2}$ in $c_{2}$ to vanish. Consequently, the error term in Eq. (83) is of order $(b/b_{\mathrm{c}}-1)\log(b/b_{\mathrm{c}}-1)$.
@@ VI Summary and discussion
Furthermore, using our results, we show that within the eikonal approximation, the QNM frequencies can be expressed in terms of the local geometric and material quantities at the photon sphere. The real part of the QNM frequencies is related to the orbital frequency at the photon sphere, while the imaginary part is determined by the Lyapunov exponent of the unstable photon circular orbit. In particular, we obtain
@@
Thus, the QNM frequencies are ultimately determined by the same local matter field (or local curvature) quantities that govern the deflection angle. [^4] It should be noted that the correspondence between photon spheres and eikonal QNMs is not strictly universal; for example, in Lovelock gravity, the eikonal QNMs frequencies of gravitational perturbations deviate from the photon-sphere prediction due to the spin-dependent structure of the effective potential [44]. Therefore, precise measurements of QNM frequencies—such as those from gravitational wave observations—provide an independent means to probe the local curvature and matter distribution near the photon sphere, thereby reinforcing the connection between strong gravitational lensing and the dynamical properties of compact objects. Furthermore, our findings may also shed light on a universal upper bound on chaos in thermal quantum field theory [45, 46], which provides an inequality between the Lyapunov exponent and the surface gravity of the horizon, a relation whose generalization to the photon sphere has recently been discussed [47].
@@ Acknowledgements.
The author gratefully acknowledges useful discussions and valuable comments from Ken-ich Nakao, Tomohiro Harada, Tsutomu Kobayashi, Yasuyuki Hatsuda, Hideki Ishihara, Chulmoon Yoo, Akio Hosoya, Yohsuke Takamori, Koji Nakamura, Masashi Kimura, Tetsuya Shiromizu, Hiroshi Kozaki, Shinya Aoki, and Naoki Tsukamoto. This work was supported in part by JSPS KAKENHI Grant Nos. JP22K03611, JP23KK0048, and JP24H00183.

### v4

@@ I Introduction
Building on these insights, numerous studies have focused on developing analytical formulations for the deflection angle in the strong deflection limit and exploring their applications. An explicit expression for the deflection angle in this context was first derived in Ref. [5], establishing the foundation for subsequent research. Later, this approach was refined in Refs. [6, 7], thus extending its applicability to various spacetimes. Because of significant astrophysical interest in this regime and the methodological simplicity of these techniques, they have been applied to a broad range of compact astrophysical objects. Notable examples include the Schwarzschild black hole [8], the Reissner–Nordström black hole [9, 10], regular black holes [11], naked singularities [12, 13], wormholes [14, 15, 16, 17, 18, 19, 20], ultracompact objects [7], gravastars [21], and other black holes [22, 23]. Additionally, formulations in the strong deflection limit have been extended to incorporate finite-distance corrections [24, 25] and to describe the deflection of massive particles [26]. Consequently, the logarithmic divergence of the deflection angle in this regime has been consistently verified across a wide range of spacetimes.
@@
Here, $b$ denotes the impact parameter, and $b_{\mathrm{c}}$ is its critical value. This formula has been successfully applied to a wide variety of astrophysical scenarios. However, two fundamental issues remain unresolved in the current formulation. First, the physical interpretation of $\bar{a}$ and $\bar{b}$ is not yet fully understood. Although these coefficients are expressed in terms of the local values of metric functions and their derivatives (and are therefore practically useful), their precise relationship to physical quantities—such as the local spacetime geometry or the influence of matter fields—remains unclear. Second, these coefficients, in their current form, are coordinate dependent. Since they are related to observable quantities, it is essential to reformulate them in a coordinate-independent manner in accordance with the principles of relativity, thus enabling a more universal physical interpretation.
@@ II Photon dynamics and circular orbits in static, spherically symmetric spacetimes
where $t$, $r$, and $(\theta,\varphi)$ denote the time, radial, and angular coordinates, respectively. The functions $\nu(r)$, $\lambda(r)$, and $R(r)$ are arbitrary; however, one of these can be fixed by an appropriate gauge choice. Here, $R(r)$ represents the areal radius [i.e., $R(r)>0$]. We assume that the spacetime is asymptotically flat, meaning that as $R\to\infty$ (corresponding to $r\to r_{\infty}$, with $r_{\infty}$ denoting the coordinate value at spatial infinity), the metric functions approach $e^{\nu}\to 1$ and $e^{\lambda}\to(R^{\prime})^{2}$, ensuring convergence to the Minkowski metric, [^1] In our formulation, we do not impose the coordinate condition that the radial coordinate asymptotically approaches the areal radius at infinity, as is commonly done in the literature (e.g., Refs. [5, 6]). where, and hereafter, the prime denotes differentiation with respect to $r$.
@@
provided that $R>2m$, which ensures that the metric is well defined.
@@
Because of the spherical symmetry of the metric, their orbits can be confined to the equatorial plane (i.e., $\theta=\pi/2$) without loss of generality. Under these assumptions, the Lagrangian for a photon is given by
@@ III Derivation of the deflection angle in the strong deflection limit
Following Ref. [5], we evaluate the deflection angle in the strong deflection limit. We introduce the new variable,
@@
which differs from the variable used in the literature. This definition ensures that the final expressions are formulated primarily in terms of the areal radius. As a result, the formulation is both coordinate invariant and geometrically meaningful. With this substitution, the integral $I(r_{0})$ becomes
@@ III.1 $R^{\prime}_{\mathrm{m}}\neq 0$
$c_{2}$ $=-\frac{V^{\prime\prime}_{\mathrm{m}}}{2}+O\left(\frac{r_{0}}{r_{\mathrm{m}}}-1\right).$ (30)
@@
While previous works in the literature have expressed these coefficients arising from $I_{\mathrm{D}}$ in terms of metric functions, our formulation instead recasts them in terms of local quantities at the photon sphere (i.e., $V^{\prime\prime}_{\mathrm{m}}$ and $R_{\mathrm{m}}$) together with a well-defined quasi-local quantity [i.e., $m(r_{\mathrm{m}})$]. This approach not only clarifies the physical origin of the logarithmic divergence in the deflection angle but also identifies the observable quantities governing the phenomenon in the strong deflection limit.
@@ III.2 $R^{\prime}_{\mathrm{m}}=0$
Next, we assume that $R^{\prime}_{\mathrm{m}}=0$, which implies that the spacetime exhibits a wormhole geometry with the photon sphere located at the throat. In this case, the expansion of $c_{1}$ and $c_{2}$ around $r_{0}=r_{\mathrm{m}}$ reduces to [^3] Although the first-order term in $(r_{0}/r_{\mathrm{m}}-1)$ in the expansion of $c_{2}$ does not affect the strong field limit coefficients, we retain it to capture the unique behavior of the higher-order correction terms for the Ellis–Bronnikov wormhole.
@@ IV Formula for the deflection angle in the strong deflection limit
$\nu^{\prime}$ $=\frac{Re^{\lambda}}{R^{\prime}}G_{(1)(1)}-\frac{(R^{\prime})^{2}-e^{\lambda}}{RR^{\prime}}.$ (53)
@@
Consequently, these results imply that information about the curvature near the photon sphere can be extracted from measurements of the deflection angle in the strong deflection regime. Specifically, if $\bar{a}$ and $R_{\mathrm{m}}$ are determined from observations, the curvature combination $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$ can be directly derived. Moreover, if $\bar{b}$ is also measured and a specific spacetime model is assumed—so that all the relevant quantities can be computed for direct comparison with observations—then we can deduce $G_{(1)(1)}^{\mathrm{m}}$ [with $m(r_{\mathrm{m}})$] determined via Eq. (55) in the case $R^{\prime}_{\mathrm{m}}\neq 0$, or $G_{(0)(0)}^{\mathrm{m}}$ (with $\mathrm{d}m/\mathrm{d}R|_{r=r_{\mathrm{m}}}$) determined via Eq. (56) in the case $R^{\prime}_{\mathrm{m}}=0$. In other words, if the theoretical predictions for $\bar{a}$ and $\bar{b}$ agree reasonably well with the observed values, this approach enables us to infer the local curvature and the gravitational mass, thereby providing a robust consistency check of the underlying gravitational model.
@@ V Interpreting strong deflection in terms of local matter fields in general relativity
Finally, we express the strong field limit coefficients $\bar{a}$ and $\bar{b}$ in terms of these matter field quantities. By substituting Eq. (62) into Eqs. (60) and (69), we obtain
@@
As the simplest case, we consider a vacuum spacetime in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Schwarzschild line element,
@@
Next, we consider an electrovacuum spacetime (i.e., a vacuum spacetime containing only electromagnetic fields) in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Reissner–Nordström line element,
@@
Next, consider a special case in which the spacetime is nonvacuum, but matter fields satisfy $\rho(r_{\mathrm{m}})+\Pi(r_{\mathrm{m}})=0$. In this situation, the strong field limit coefficient $\bar{a}$ remains equal to $1$, thereby mirroring the Schwarzschild result despite the presence of nonvanishing matter fields. This demonstrates that, when $\rho(r_{\mathrm{m}})+\Pi(r_{\mathrm{m}})=0$, the strong deflection limit exhibits universal behavior with $\bar{a}=1$, irrespective of the specific details of the matter content.
@@
which agree with those in Ref. [14], employing the same expansion parameter used here. Note that for the Ellis–Bronnikov wormhole under consideration, the conditions $R^{\prime\prime\prime}_{\mathrm{m}}=0$ and $V^{\prime\prime\prime}_{\mathrm{m}}=0$ are satisfied, which causes the coefficient of $(b/b_{\mathrm{c}}-1)^{1/2}$ in $c_{2}$ to vanish. Consequently, the error term in Eq. (83) is of order $(b/b_{\mathrm{c}}-1)\log(b/b_{\mathrm{c}}-1)$.
@@ VI Summary and discussion
Furthermore, using our results, we show that, within the eikonal approximation, the QNM frequencies can be expressed in terms of the local geometric and material quantities at the photon sphere. The real part of the QNM frequencies is related to the orbital frequency at the photon sphere, while the imaginary part is determined by the Lyapunov exponent of the unstable photon circular orbit. In particular, we obtain
@@
Thus, the QNM frequencies are ultimately determined by the same local matter field (or local curvature) quantities that govern the deflection angle. [^4] It should be noted that the correspondence between photon spheres and eikonal QNMs is not strictly universal; for example, in Lovelock gravity, the eikonal QNM frequencies of gravitational perturbations deviate from the photon-sphere prediction due to the spin-dependent structure of the effective potential [43]. Therefore, precise measurements of QNM frequencies—such as those from gravitational wave observations—provide an independent means to probe the local curvature and matter distribution near the photon sphere, thereby reinforcing the connection between strong gravitational lensing and the dynamical properties of compact objects. Furthermore, our findings may also shed light on a universal upper bound on chaos in thermal quantum field theory [44, 45], which provides an inequality between the Lyapunov exponent and the surface gravity of the horizon, a relation whose generalization to the photon sphere has recently been discussed [46].
@@ Acknowledgements.
The author gratefully acknowledges useful discussions and valuable comments from Ken-ich Nakao, Tomohiro Harada, Tsutomu Kobayashi, Yasuyuki Hatsuda, Hideki Ishihara, Chulmoon Yoo, Akio Hosoya, Yohsuke Takamori, Koji Nakamura, Masashi Kimura, Tetsuya Shiromizu, Hiroshi Kozaki, Shinya Aoki, and Naoki Tsukamoto. This work was supported in part by JSPS KAKENHI Grants No. JP22K03611, No. JP23KK0048, and No. JP24H00183 and by Gakushuin University.